### v1

@@ 5.1 Synchronous Deployment
Latency-sensitive use cases are sometimes deployed in synchronous mode, i.e., only one query is processed at a time. Synchronous deployment leads to the lowest latency since different requests do not compete for computation, and it tends to be dominated by decode time.
Table 5 compares the inference performance for synchronous deployment across different model sizes, GPU types, quantization schemes, and use cases. For each configuration, we highlight the number of GPUs that are most cost-effective. The results reveal that W4A16-INT quantization consistently delivers the highest performance gains for synchronous deployment across all model sizes, GPU types, and latency-sensitive use cases. For both the 8B and 70B model sizes, W4A16-INT achieves significant 2-3x cost reduction per query and 1.5-2x latency improvements compared to the baseline full-precision BF16 format.
For the 405B model, W4A16-INT is exceptionally effective, offering 6–7x reductions in cost per query and enabling inference on a smaller number of GPUs. Specifically, deploying the 405B model with W4A16-INT on 4x A100 or H100 GPUs meets performance thresholds that previously required 16 GPUs in BF16 format, reducing inter-GPU communication and lowering latency. As seen in the previous section, the accuracy differences are minor.
@@ 5.2 Asynchronous Deployment
Processing multiple queries simultaneously generally results in better computational efficiency than a single query at a time. vLLM automatically coordinates multiple asynchronous queries, balancing computation between prefill and decode stages of parallel requests. Asynchronous deployment usually results in increased latency compared to synchronous deployment, but the processing time is amortized across multiple queries, leading to significant gains in average queries per second. For latency-sensitive use cases, we report the maximum inference throughput (queries per second) under the constraints listed in Table 4, whereas for “throughput” use cases, we consider the absolute maximum inference throughput.
Table 6 shows the inference performance for the asynchronous deployment of both latency-sensitive and throughput use cases. The hardware configurations were optimized to match the number of GPUs yielding the best performance for full-precision BF16 deployment, and we assume optimal orchestration for the quantized models.
Mdl. HW Fmt. Speedup Code Compl. Doc. Gen. Code Fixing RAG Inst. Following Multi-Turn Chat Sum. (S) Sum. (L)
QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$
8B 1×A6000 BF16 4.5 20.2k 5.5 24.8k 2.2 10.1k 4.2 18.9k 10.3 46.5k 5.3 23.7k 4.3 19.5k 1.0 4.3k
INT8 1.16 4.5 20.3k 7.0 31.4k 1.9 8.7k 5.2 23.3k 16.6 74.5k 5.1 22.9k 5.8 26.3k 1.1 5.1k
INT4 1.56 8.0 35.8k 9.5 42.8k 3.9 17.7k 7.4 33.4k 17.8 80.2k 8.7 39.2k 4.0 17.9k 0.9 3.9k
70B 4×A6000 BF16 0.7 794 0.4 491 0.4 478 0.6 631 1.3 1.4k 0.5 537 1.4 1.6k 0.3 312
INT8 3.10 2.4 2.7k 1.5 1.6k 1.2 1.4k 1.4 1.5k 3.6 4.1k 1.8 2.0k 2.5 2.8k 0.4 504
INT4 3.17 3.1 3.5k 1.6 1.8k 0.8 947 0.7 814 3.8 4.3k 2.2 2.5k 3.1 3.5k 0.3 363
4×A100 BF16 5.5 2.8k 3.1 1.6k 2.3 1.2k 2.6 1.3k 6.2 3.1k 3.6 1.8k 4.2 2.1k 1.0 514
INT8 1.75 7.6 3.8k 5.7 2.9k 4.1 2.0k 4.7 2.4k 11.6 5.8k 6.7 3.4k 6.9 3.5k 1.7 844
INT4 1.33 7.7 3.9k 4.0 2.0k 2.7 1.4k 3.0 1.5k 9.5 4.8k 5.2 2.6k 7.7 3.9k 0.9 443
4×H100 BF16 11.7 3.4k 7.6 2.2k 5.4 1.6k 6.4 1.9k 13.3 3.9k 7.9 2.3k 8.8 2.6k 2.2 630
FP8 1.67 19.4 5.6k 13.0 3.8k 8.9 2.6k 10.5 3.1k 22.9 6.7k 12.7 3.7k 19.4 5.6k 4.3 1.3k
INT4 1.10 13.5 3.9k 7.9 2.3k 5.4 1.6k 5.9 1.7k 17.1 5.0k 9.5 2.8k 13.5 4.0k 1.6 466
405B 16×A100 BF16 0.4 48 0.4 47 0.2 25 0.3 38 0.8 96 0.3 42 2.2 276 0.4 56
INT8 8.09 3.7 461 2.1 269 1.8 228 2.1 270 6.0 755 3.1 390 3.7 470 0.9 119
INT4 7.12 3.7 464 2.2 280 1.5 190 1.6 202 5.1 639 2.5 320 2.8 353 0.6 70
16×H100 BF16 1.8 137 1.4 109 0.8 63 1.2 92 2.7 204 1.4 107 4.0 302 0.8 62
FP8 5.61 8.3 627 8.2 618 4.9 366 7.1 533 15.1 1.1k 8.8 663 6.2 464 1.8 133
INT4 3.53 7.2 542 4.4 330 3.1 235 3.4 257 10.1 757 5.6 423 4.6 344 1.0 77
QPS: Queries per second (higher is better). Q/\$: Queries per USD (higher is better). Sum. (S/L): Small/Large Summarization.
Numbers denoted with k represent thousands (e.g., 20.3k = 20,300).
The data indicates that W8A8-INT and W8A8-FP are the most effective quantization schemes for the asynchronous deployment of mid- and large-size models on high-end GPUs. As expected, W8A8-FP performs best on H100 GPUs, while W8A8-INT is more suited for efficient inference on A100 GPUs.
In contrast, W4A16-INT is the best option for cost efficiency on A6000 GPUs. It outperforms other quantization schemes, especially for smaller models and latency-sensitive tasks. It suggests that W4A16-INT is a viable and cost-effective option for less memory-intensive applications on mid-tier hardware. When evaluating cost efficiency across different GPU classes for mid- and large-size models, W8A8-FP on H100 GPUs consistently stands out as the best option across all use cases.
Overall, the results underscore no universally optimal quantization scheme for asynchronous deployment. The best choice depends on model size, GPU type, and whether the task prioritizes latency or throughput. Nonetheless, quantization leads to significant cost reduction per query regardless of the scenario, achieving 1.5-3x cost reduction for 8B and 70B and 5-8x for 405B models.
@@ 6 Conclusion
We provided a broad and in-depth study of accuracy-vs-performance-vs-cost trade-offs for quantized LLMs across various deployment environments. Specifically, we covered all quantization formats with efficient computational support and a range of quantization algorithms, deployment use cases, and GPUs. In Figure 7 we summarize our findings in terms of accuracy recovery per quantization format, using carefully-tuned state-of-the-art quantization techniques.
Broadly speaking, our findings show that, with a judicious choice of algorithm and parametrization, these formats can offer higher accuracy than previously thought, significantly improve inference performance, and reduce costs. At the same time, we have also shown that the optimal choice of format can be task and algorithm specific, providing guidelines for this choice.
Acknowledgements
We would like to express our sincere gratitude to the Neural Magic team, the vLLM community, and EleutherAI for developing the libraries that enabled this research. Specifically, we used the LLM Compressor to produce most of our quantized models, vLLM as an inference engine to accelerate evaluations, and the LM Evaluation Harness to run all of our academic benchmarks.
@@ A.1 Real-World Benchmarks
In Tables 8 and 9 we report pass@10 scores for all models on HumanEval and HumanEval+ benchmarks.
@@ A.4 Academic Benchmarks
In Tables 10 and 11 we report accuracy recoveries per-task across academic benchmarks.

### v2

@@ 5.1 Synchronous Deployment
Latency-sensitive applications are sometimes deployed in synchronous mode, where a single query is processed at a time. This approach minimizes latency by avoiding resource contention, making inference largely decode-bound.
Table 4 compares inference performance across model sizes, GPU types, quantization schemes, and use cases, highlighting the most cost-effective GPU configurations. The results show that W4A16-INT consistently achieves the highest performance gains across all models and hardware setups.
For 8B and 70B models, W4A16-INT reduces cost per query by 2–3× and improves latency by 1.5–2.5× compared to the full-precision BF16 baseline. The impact is even more pronounced at 405B, where W4A16-INT achieves 5–7× cost reductions and enables inference with fewer GPUs. Notably, deploying the 405B model on 4× A100 or H100 GPUs with W4A16-INT meets performance thresholds that previously required 16 GPUs in BF16, reducing inter-GPU communication and latency. Given the minor accuracy trade-offs observed in the previous section, this makes W4A16-INT highly effective for synchronous deployment.
@@ 5.2 Asynchronous Deployment
Processing multiple queries concurrently improves computational efficiency compared to single-query execution. vLLM automatically manages asynchronous requests, balancing computation between prefill and decode stages.
While asynchronous deployment increases per-query latency relative to synchronous execution, it amortizes computation across multiple requests, significantly boosting overall throughput, measured in queries per second (QPS). Table 5 reports the maximum achievable throughput and cost efficiency (queries per dollar) across different quantization formats, model sizes, and hardware configurations. The setups were optimized for peak BF16 performance and kept consistent when evaluating quantized models. Results show that W8A8-INT and W8A8-FP yield the highest throughput, though W4A16-INT remains competitive and can outperform W8A8 in some scenarios.
Many real-world applications impose latency constraints on asynchronous deployment. Figures 2 and  3 illustrate trade-offs between latency and throughput for two example tasks: docstring generation and code fixing. W4A16-INT is more efficient at lower latencies, making it ideal for applications requiring rapid response times. In contrast, W8A8 formats maximize throughput at the cost of higher latency, making them better suited for batch processing. The point where W8A8 overtakes W4A16 depends on factors such as model size, hardware, and task requirements.
Figure 2: Latency-throughput example for docstring generation use-case. W4A16 is more efficient at low latency (lower throughput), whereas W8A8 becomes more efficient at high latency (high throughput).
Figure 3: Latency-throughput example for code fixing use-case. W4A16 is more efficient at low latency (lower throughput), whereas W8A8 becomes more efficient at high latency (high throughput).
@@ 6 Conclusion
We provided a broad, in-depth study of accuracy-vs-performance-vs-cost trade-offs for quantized LLMs across various deployment environments, covering all quantization formats with efficient support, and a range of quantization algorithms, deployment use cases, and GPUs. In Figure 4 we summarize our findings in terms of accuracy recovery per quantization format, using carefully-tuned state-of-the-art quantization techniques. Broadly, our findings show that, with a judicious choice of algorithm and parametrization, these formats can offer higher accuracy than previously thought, significantly improve inference performance, and reduce costs. At the same time, we have also shown that the optimal choice of format can be task and algorithm specific, providing guidelines for this choice.
Limitations
While our study provides a comprehensive evaluation of quantization effects on model accuracy and inference performance, several limitations remain. We have primarily focused on weight and activation quantization, leaving open questions about the impact of compressing other model components such as the KV-cache, input embeddings, and language modeling head. Further investigation is needed to assess how these additional compression techniques influence both accuracy and computational efficiency. Additionally, our analysis does not fully explore the effects of quantization across specialized use cases, such as multi-lingual tasks, where accuracy degradation could vary significantly depending on the language distribution and underlying model architecture. Future work should extend these evaluations to provide a more holistic understanding of quantization trade-offs in diverse deployment scenarios.
@@ A.1 Real-World Benchmarks
In Figures 5 and 6 we report pass@10 scores for all models on HumanEval and HumanEval+ benchmarks.
@@ A.4 Academic Benchmarks
In Tables 9 and 10 we report accuracy recoveries per-task across academic benchmarks.